Nadia can Completely mow the yard by herself in 30 minutes. Peter can completely mow the yard by himself in 45 minutes. How long would it take both of them to mow the yard if they worked together?
Answer by mananth(16946)

Nadia can Completely mow the yard by herself in 30 minutes. Peter can completely mow the yard by himself in 45 minutes. How long would it take both of them to mow the yard if they worked together?
...
Nadia 30 minutes
Peter = 45 minutes
..
Nadia = 1/30 of job in 1 minute
Peter = 1/45 of the job in1 minute.
Together = 1/30 +1/45
LCD = 90
=(3+2)/90
=5/90
=1/18 of the job in 1 minute.
together they will take 18 minutes to complete the job.
,,,.
m.ananth@hotmail.ca